OpenStudy (anonymous):

0.12(x-10)-0.05(x-30)=-1.80 how do i solve this?

OpenStudy (anonymous):

Do the multplications: 0.12x - 0.12*10 - 0.05x + 0.05*30 = -1.80

OpenStudy (anonymous):

Combine like terms

OpenStudy (anonymous):

1.2-1.5?

OpenStudy (anonymous):

that what i got?

OpenStudy (anonymous):

i need help!

OpenStudy (anonymous):

Change the decimal fractions to fractions:\[\frac{12}{100}(x-10)-\frac{5}{100}(x-30)=\frac{-18}{10} \]Expand the products.\[-\frac{6}{5}+\frac{3 x}{25}+\frac{3}{2}-\frac{x}{20}=\frac{-18}{10} \]Collect like tems on the LHS.\[\frac{3}{10}+\frac{7 x}{100}=\frac{-18}{10} \]Move the 3/10 to the RHS.\[\frac{7 x}{100}=\frac{-18}{10}-\frac{3}{10} \]Collect the terms in the RHS.\[\frac{7 x}{100}=-\frac{21}{10} \]Multiply each side by 100/7.\[\frac{100}{7}\frac{7 x}{100}=-\frac{21}{10}\frac{100}{7} \]Simplify the result.\[x=-30 \]

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

thank you for ur help!